OpenStudy (minatsukisaya95):

A small funnel in the shape of a cone is being emptied of fluid at the rate of 12 cm cube per second. The height of the funnel is 20 cm and the radius of the top is 4 cm. How fast is the fluid level dropping when the level stands 5 cm above the vertex of the cone?

OpenStudy (anonymous):

let r be the radius and h be the height at any instant then r/h = 4/20 or r/h =1/5 or r=(1/5) h...

OpenStudy (anonymous):

now volume at any instant is given by V=(pi)*(r^2)*h or V=(pi)*((h/5)^2)*h V=((pi)* h^3)/25 V=((pi)/25)*h^3 now differentiating wrt t both sides we have dV/dt =3*(pi/25)*h^2 *(dh/dt) now given dV/dt =-12 and h=5 ...we need (dh/dt) just solve it to get (dh/dt)

OpenStudy (minatsukisaya95):

thank you i got it

OpenStudy (anonymous):

yw

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."